OpenStudy (anonymous):

|x| = 4 |x| − 8 = −5 −3|2x + 6| = −12

OpenStudy (anonymous):

Are there 3 separate questions?

OpenStudy (anonymous):

For the first one x=4 or x=-4

OpenStudy (anonymous):

yes marinda

OpenStudy (anonymous):

Do you know how to solve absolute equations, do you need help?

OpenStudy (anonymous):

ALOT of help... i kinda had a dumba** 4 a teacher

OpenStudy (anonymous):

That's fine, i'm sure you're not dumb. Solving inequalities isn't that hard. Instead of solving one equation you have to solve two. The positive answer and the negative one. Let's look at |x| − 8 = −5 First step is to get the absolute part by itself, can you do that?

OpenStudy (anonymous):

sorry i meant to write solving absolutes :)

OpenStudy (anonymous):

i have no idea how to do that.

OpenStudy (anonymous):

That okays so |x| − 8 = −5 We want to take the 8 over to get the absolute by itself, to do this add 8 to both sides |x|-8+8=-5+8 |x|=3

OpenStudy (anonymous):

so basically its just like solving a normal equation

OpenStudy (anonymous):

exactly here's where it differs. because an absolute always gives a positive value x=3 or x=-3

OpenStudy (anonymous):

so your always going to have two answers one positive and one negative?

OpenStudy (anonymous):

yes but they are not necessarily the same number. let's try the last one it will help explain −3|2x + 6| = −12 now make the absolute part on it's own. what do you get?

OpenStudy (anonymous):

are you able to do that? it's just like solving for a normal equation

OpenStudy (anonymous):

|2x+6|=4 So 2x+6=4 or 2x+6=-4 Solve like normal, hope it helps

OpenStudy (anonymous):

i get it now.. thank you very much